# Neurons: Emulating the Human Visual Cortex Improves Fidelity and Interpretability in fMRI-to-Video Reconstruction

### v1

@@ Abstract
Decoding visual stimuli from neural activity is essential for understanding the human brain. While fMRI methods have successfully reconstructed static images, fMRI-to-video reconstruction faces challenges due to the need for capturing spatiotemporal dynamics like motion and scene transitions. Recent approaches have improved semantic and perceptual alignment but struggle to integrate coarse fMRI data with detailed visual features. Inspired by the hierarchical organization of the visual system, we propose NEURONS, a novel framework that decouples learning into four correlated sub-tasks: key object segmentation, concept recognition, scene description, and blurry video reconstruction. This approach simulates the visual cortex’s functional specialization, allowing the model to capture diverse video content. In the inference stage, NEURONS generates robust conditioning signals for a pre-trained text-to-video diffusion model to reconstruct the videos. Extensive experiments demonstrate that NEURONS outperforms state-of-the-art baselines, achieving solid improvements in video consistency (26.6%) and semantic-level accuracy (19.1%). Notably, NEURONS shows a strong functional correlation with the visual cortex, highlighting its potential for brain-computer interfaces and clinical applications. Code and model weights will be available at: https://github.com/xmed-lab/NEURONS.
@@ 1 Introduction
We believe emulating the functional architecture of the visual cortex can alleviate the issue and enhance the decoding of brain activity. Studies [24, 16, 21, 43] have shown that different regions of the visual cortex process varying levels of visual information granularity (as depicted in Fig. 1). Drawing inspiration from the hierarchical structure of the human visual system, we decompose the learning process into multiple sub-tasks, enabling the brain model to capture diverse aspects of video content. Specifically, early visual areas such as V1 and V2 are responsible for encoding low-level features like edges and shapes, akin to the segmentation task in computer vision. However, segmenting every object in a video is impractical due to the inherent limitations of human memory in retaining detailed video information. Since individuals predominantly focus on key objects and salient features within scenes rather than processing all finer details [17], we introduce the key object segmentation task, which trains the brain model to initially learn the shapes and contours of key objects. Subsequently, we progressively incorporate more complex semantic concepts and motion information through the concept recognition task, which mimics the object and face recognition processes in areas V4 and ITC, compensating for the semantic loss incurred by focusing solely on key objects. Then, the scene description task enables the brain model to identify correlations between concepts and generate coherent scene descriptions. Finally, the blurry video reconstruction task integrates motion and color distribution information. These four tasks are learned using a progressive strategy that mirrors the hierarchical structure of the visual cortex.
@@ 3.2 Neurons - Brain Model
The brain model aims to encode fMRI voxels into video embeddings. Its training serves as a pre-training stage to produce suitable video embeddings, preparing for subsequent decoupled tasks. After preprocessing, a video is split into several clips at two-second intervals (i.e., fMRI time resolution), each video clip $y_{c}$ has six frames $y_{c}\in\mathbb{R}^{B\times F\times C\times H\times W}$ where $F$ the number of frames and $F=6$. $x_{c}$ is the corresponding fMRI signal of video clip $c$. We use the same pre-trained MindEyeV2 [33] (details are illustrated in Supplementary §B.1). Previous approaches like NeuroClips directly align the image embeddings from the backbone, overlooking temporal relationships between consecutive frames. Instead, we use a motion projection $\mathcal{P}_{vid}(\cdot)$ to map the image embeddings $e^{i}\in\mathbb{R}^{B\times N\times C}$ to spatial-temporal space $e^{v}\in\mathbb{R}^{B\times F\times N\times C}$. We use the same bidirectional contrastive loss as in [33, 10] called BiMixCo, denoted as $\mathcal{L}_{\mathrm{CLIP_{v}}}$, which combines MixCo and contrastive loss, to align $e_{v}$ with the target video embeddings of CLIP vision encoder, detailed in Supplementary §B.1. Then, we embed $e^{v}$ to obtain the text embedding $e^{t}$. A contrastive learning loss between CLIP text embedding and $e^{t}$ is adopted to train the additional text modality. The contrastive loss serves as the training loss $\mathcal{L}_{\mathrm{CLIP_{t}}}$ of this process, similar to $\mathcal{L}_{\mathrm{CLIP_{v}}}$, omitted here. This pre-alignment provides feasible vision and text features which are then fed into the Decoupler for decoupled task learning.
@@ 3.3 Neurons - Decoupler: Decoupled Explicit Tasks Learning
$\mathcal{L}_{cls}=\mathcal{L}_{ce}(\mathcal{D}_{cls}(\bar{e^{v}}),\mathcal{C})$ (1)
@@
$\mathcal{L}_{txt}=-\frac{1}{|\mathcal{S}|}\sum^{|\mathcal{S}|}_{i=1}\log\mathcal{D}_{txt}(s_{i}|s_{<i},e^{t})$ (2)
@@ 4.1 Experimental Setup
Table 1: Quantitative comparison of Neurons reconstruction performance against other methods (Video-based). Bold font signifies the best performance, while underlined text indicates the second-best performance. MinD-Video and NeuroClips are both results averaged across all three subjects, and the other methods are results from subject 1. Results of baselines are quoted from [10].
@@ 4.2 Comparison with State-of-the-Art Methods
In this section, we compare Neurons with previous video reconstruction methods on the cc2017 dataset both quantitatively and qualitatively. In quantitative metrics, our proposed method demonstrates superior performance across all video-based metrics compared to existing methods, as shown in Table 1. Semantic-level 2-way accuracy achieves 0.863, surpassing the previous best result (0.839 by MinD-Video) by a relative margin of 2.9%. Notably, our method significantly outperforms state-of-the-art methods in ST-level CLIP-pcc, attaining 0.934, which represents a 26.6% improvement over NeuroClips (0.738). Furthermore, our method achieves the highest 0.262 in the 50-way accuracy, outperforming NeuroClips (0.220) by 19.1%. The consistent superiority across all metrics, coupled with lower standard deviations, underscores the robustness of our approach. These results validate that our method effectively bridges semantic and spatiotemporal feature learning, setting a new benchmark for video understanding tasks. Moreover, we also provided a frame-based comparison, which is detailed in Supplementary §E.1 and Table a. For qualitative comparison, as can be seen in Fig. 5, compared to NeuroClips, our Neurons generates videos with higher quality, more precise semantics (e.g., people, turtles, and airplanes), and smoother movements. Furthermore, thanks to the decoupled tasks, Neurons has a better capability of reconstructing more complex scenes, such as a crowd of people. More visualizations and failure case analysis can be found in Supplementary §E.2.
@@ 5 Explicit Evaluation of Decoupled Tasks
Scene Description. The previous method, NeuroClips, generates text prompts for the T2V diffusion model by captioning the generated keyframes using BLIP. However, this approach lacks robustness due to inaccuracies in the generated keyframes, which often hinder the precise identification of semantic content, as demonstrated in Table 3. We conduct a comprehensive comparison between the descriptions generated by NeuroClips and our method. The results reveal that Neurons produces more accurate scene descriptions at the sentence level, particularly in capturing verbs that reflect better motion information. Further details can be found in Supplementary §E.5.
@@ 6 Interpretation Results
Inspired by the human visual cortex, we designed our Neurons with four different decoupled tasks. In order to validate our initial inspiration, we map the projected embeddings of each decoupled task to a brain map using BrainDecodesDeepNets [46], a visualization tool that uncovers functional correlations between Neurons’s tasks and distinct regions of the visual cortex. As shown in Figure 6, the embeddings align with specific regions: $e^{seg}$ with V1, V2, and MT (due to video-based segmentation), $e^{cls}$ with V4, EBA and LOC, $e^{txt}$ with PPA and FFA. Notably, $e^{rec}$ highlights V3 areas while also encoding richer semantic information. This alignment with various visual cortex regions validates our design principles and underscores the biological plausibility of our approach. More details are in Supplementary §F.

### v2

@@ Abstract
Decoding visual stimuli from neural activity is essential for understanding the human brain. While fMRI methods have successfully reconstructed static images, fMRI-to-video reconstruction faces challenges due to the need for capturing spatiotemporal dynamics like motion and scene transitions. Recent approaches have improved semantic and perceptual alignment but struggle to integrate coarse fMRI data with detailed visual features. Inspired by the hierarchical organization of the visual system, we propose NEURONS, a novel framework that decouples learning into four correlated sub-tasks: key object segmentation, concept recognition, scene description, and blurry video reconstruction. This approach simulates the visual cortex’s functional specialization, allowing the model to capture diverse video content. In the inference stage, NEURONS generates robust conditioning signals for a pre-trained text-to-video diffusion model to reconstruct the videos. Extensive experiments demonstrate that NEURONS outperforms state-of-the-art baselines, achieving solid improvements in video consistency (26.6%) and semantic-level accuracy (19.1%). Notably, NEURONS shows a strong functional correlation with the visual cortex, highlighting its potential for brain-computer interfaces and clinical applications. Code and model weights are available at: https://github.com/xmed-lab/NEURONS.
@@ 1 Introduction
Emulating the functional architecture of the visual cortex offers a promising strategy to improve the decoding of brain activity. Prior research has demonstrated that distinct regions within the visual cortex are specialized for processing different levels of visual information granularity [25, 16, 21, 44] (see Fig. 1). Inspired by the hierarchical organization of the human visual system, we design the learning framework as a composition of multiple sub-tasks, each reflecting specific perceptual roles. For example, early visual areas such as V1 and V2 are involved in encoding low-level visual features—such as edges and shapes—which parallels the function of segmentation modules in computer vision. However, segmenting every object in a video is impractical due to the inherent limitations of human memory in retaining detailed video information. Since individuals predominantly focus on key objects and salient features within scenes rather than processing all finer details [17], we introduce the key object segmentation task, which trains the brain model to initially learn the shapes and contours of key objects. Subsequently, we progressively incorporate more complex semantic concepts and motion information through the concept recognition task, which mimics the object and face recognition processes in areas V4 and ITC, compensating for the semantic loss incurred by focusing solely on key objects. Then, the scene description task enables the brain model to identify correlations between concepts and generate coherent scene descriptions. Finally, the blurry video reconstruction task integrates motion and color distribution information. These four tasks are learned using a progressive strategy that mirrors the hierarchical structure of the visual cortex.
@@ 3.2 Neurons - Brain Model
The brain model aims to encode fMRI voxels into video embeddings. Its training serves as a pre-training stage to produce suitable video embeddings, preparing for subsequent decoupled tasks. After preprocessing, a video is split into several clips at two-second intervals (i.e., fMRI time resolution), each video clip $y_{c}$ has six frames $y_{c}\in\mathbb{R}^{B\times F\times C\times H\times W}$ where $F$ the number of frames and $F=6$. $x_{c}$ is the corresponding fMRI signal of video clip $c$. We use the same pre-trained MindEyeV2 [34] (details are illustrated in Supplementary §B.1). Previous approaches like NeuroClips directly align the image embeddings from the backbone, overlooking temporal relationships between consecutive frames. Instead, we use a motion projection $\mathcal{P}_{vid}(\cdot)$ to map the image embeddings $e^{i}\in\mathbb{R}^{B\times N\times C}$ to spatial-temporal space $e^{v}\in\mathbb{R}^{B\times F\times N\times C}$. We use the same bidirectional contrastive loss as in [34, 10] called BiMixCo, denoted as $\mathcal{L}_{\mathrm{CLIP_{v}}}$, which combines MixCo and contrastive loss, to align $e^{v}$ with the target video embeddings of CLIP vision encoder, detailed in Supplementary §B.1. Then, we embed $e^{v}$ to obtain the text embedding $e^{t}$. A contrastive learning loss between CLIP text embedding and $e^{t}$ is adopted to train the additional text modality. The contrastive loss serves as the training loss $\mathcal{L}_{\mathrm{CLIP_{t}}}$ of this process, similar to $\mathcal{L}_{\mathrm{CLIP_{v}}}$, omitted here. This pre-alignment provides feasible vision and text features which are then fed into the Decoupler for decoupled task learning.
@@ 3.3 Neurons - Decoupler: Decoupled Explicit Tasks Learning
$\mathcal{L}_{cls}=\mathcal{L}_{ce}(\mathcal{D}_{cls}(\bar{e^{v}}),\mathcal{C}),$ (1)
@@
$\mathcal{L}_{txt}=-\frac{1}{|\mathcal{S}|}\sum^{|\mathcal{S}|}_{i=1}\log\mathcal{D}_{txt}(s_{i}|s_{<i},e^{t}),$ (2)
@@ 4.1 Experimental Setup
Table 1: Quantitative comparison of Neurons reconstruction performance against other methods (Video-based). Bold font signifies the best performance, while underlined text indicates the second-best performance. MinD-Video and NeuroClips are both results averaged across all three subjects, and the other methods are results from subject 1. Results of baselines are quoted from [23, 10].
@@ 4.2 Comparison with State-of-the-Art Methods
In this section, we compare Neurons with previous video reconstruction methods on the cc2017 dataset both quantitatively and qualitatively. In quantitative metrics, our proposed method demonstrates superior performance across all video-based metrics compared to existing methods, as shown in Table 1. Semantic-level 2-way accuracy achieves 0.863, surpassing the previous best result (0.839 by MinD-Video) by a relative margin of 2.9%. Notably, our method significantly outperforms state-of-the-art methods in ST-level CLIP-pcc, attaining 0.934, which represents a 26.6% improvement over NeuroClips (0.738). Furthermore, our method achieves the highest 0.262 in the 50-way accuracy, outperforming NeuroClips (0.220) by 19.1%. The consistent superiority across all metrics, coupled with lower standard deviations, underscores the robustness of our approach. These results validate that our method effectively bridges semantic and spatiotemporal feature learning, setting a new benchmark for video understanding tasks. Moreover, we also provided a frame-based comparison, which is detailed in Supplementary §E.1 and Table a. For qualitative comparison, as can be seen in Fig. 5, compared to NeuroClips, our Neurons generates videos with higher quality, more precise semantics (e.g., people, turtles, and airplanes), and smoother movements. Furthermore, thanks to the decoupled tasks, Neurons has a better capability of reconstructing more complex scenes, such as a crowd of people. More visualizations can be found in Supplementary §E.2.
@@ 5 Explicit Evaluation of Decoupled Tasks
Scene Description. NeuroClips [10] generates text prompts for the T2V diffusion model by captioning the generated keyframes using BLIP. However, this approach lacks robustness due to inaccuracies in the generated keyframes, which often hinder the precise identification of semantic content, as demonstrated in Table 3. We conduct a comprehensive comparison between the descriptions generated by NeuroClips and our method. The results reveal that Neurons produces more accurate scene descriptions at the sentence level, particularly in capturing verbs that reflect better motion information. Further details in Supplementary §E.5.
@@ 6 Interpretation Results
Inspired by the human visual cortex, we designed our Neurons with four different decoupled tasks. In order to validate our initial inspiration, we map the projected embeddings of each decoupled task to a brain map with [47], a visualization tool that uncovers functional correlations between Neurons’s tasks and distinct regions of the visual cortex. As shown in Figure 6, the embeddings align with specific regions: $e^{seg}$ with V1, V2, and MT (due to video-based segmentation), $e^{cls}$ with V4, EBA and LOC, $e^{txt}$ with PPA and FFA. Notably, $e^{rec}$ highlights V3 areas while also encoding richer semantic information. This alignment with various visual cortex regions validates our design principles and underscores the biological plausibility of our approach. More details are in Supplementary §F.
@@ F Details of Brain Decoding Interpretation
G Limitations and Future Work
While Neurons demonstrates strong performance in fMRI-to-video reconstruction, several limitations remain. First, the model is evaluated on a single dataset with limited subject diversity, restricting generalizability. Second, the reconstructed videos, though temporally smooth, remain low in resolution and frame rate due to the coarse temporal granularity of fMRI data. Additionally, current evaluation metrics may not fully capture perceptual realism or narrative coherence. Lastly, although the model shows functional alignment with visual cortex regions, the neurobiological interpretation remains correlational rather than causal.
@@ G Limitations and Future Work
H Overview of AI Methodologies for Neuroscience Readers
This study leverages several state-of-the-art AI models to decode and reconstruct visual experiences from brain activity. At the core is a neural network-based brain model trained using contrastive learning, which aligns fMRI signals with visual and textual embeddings from CLIP, a widely used vision-language model. The decoding process is decomposed into four explicit sub-tasks—segmentation, classification, captioning, and reconstruction—each modeled by deep learning modules optimized with task-specific loss functions. For video synthesis, we use diffusion models, a class of generative models that produce high-quality video frames conditioned on the outputs of the brain model. These AI components are organized in a biologically inspired, hierarchical manner to simulate the functional specialization of the human visual cortex.